# PointAD+: Learning Hierarchical Representations for ZS 3D Anomaly Detection

### v4

@@ 4.3.5 Complexity analysis
(a) The effect of $\alpha$ to the segmentation performance.
5 Ablation Study
@@ E Point detection Multimodal detection
5.1 Ablation on modules and key hyperparameters
Module ablation. We evaluate the effectiveness of the proposed techniques by progressively incorporating the introduced modules. As shown in Table VII, the baseline vanilla model, which represents CLIP + rendering, performs poorly on both 3D and multimodal 3D anomaly detection due to CLIP’s focus on aligning 2D object semantics rather than anomaly semantics. Adding the 3D global branch incorporates global anomaly semantics into PointAD+, improving both local and global anomaly detection performance. Although the 3D local branch further enhances performance, local accuracy on multimodal 3D detection declines due to the absence of 2D constraints, leading to inefficient multimodal fusion when integrating 2D RGB information in a plug-and-play manner. The inclusion of the 2D global branch explicitly introduces 2D anomaly information, enabling PointAD+ to achieve overall performance gains and better multimodal fusion. By focusing on 2D anomaly regions, PointAD+ develops a deeper understanding of point clouds from 2D representations, improving multimodal segmentation performance from 92.9% to 97.2% P-AUROC and from 84.4% to 90.2% AUPRO. So far, PointAD+ does not explicitly account for point relationships within point clouds. To address this, explicit point learning is introduced to model abnormal spatial relationships. PointAD+ leverages geometric characteristics, achieving notable improvements in segmentation performance. Specifically, P-AUROC increases from 95.5% to 96.3% and AUPRO rises from 84.4% to 87.5% on ZS 3D detection. On multimodal 3D detection, P-AUROC improves from 97.2% to 97.9% and AUPRO increases from 90.2% to 92.5%. Finally, cross-alignment is introduced to enhance interactions between hierarchical representations, leading to further performance gains over these metrics.
@@ 5.1 Ablation on modules and key hyperparameters
Ablation on the number of neighbors The number of neighbors $k$ determines the receptive field of each point within its local neighborhood. As illustrated in Table VIII, we further explore $k\in\{5,10,15,20,25\}$ to study its impact. For point detection, the best local accuracy is achieved at $k=10$ (96.9 AUROC and 88.7 AUPRO), and the best global accuracy is also observed at $k=10$ (82.5 AUROC and 94.8 AP). For multimodal detection, both $k=10$ and $k=20$ yield superior performance. Overall, the detection results remain stable across different values of $k$. This indicates that PointAD+ is robust to the number of neighboring points..
@@
Ablation on rendering quality. In our experiments, we relied on the Open3D Library to generate point cloud renderings, though it does not provide an API for explicitly controlling rendering quality. we evaluated PointAD on MVTec3D-AD under blur levels $\sigma\in\{1,5,9\}$. As summarized in Table XIV, the detection performance of PointAD+ decreases as the rendering quality degrades (i.e., with larger $\sigma$ values). Nevertheless, even under heavy blur ($\sigma=9$), the model maintains acceptable accuracy and consistently outperforms baselines that rely on high-quality renderings.
Ablation on point occlusion. We next evaluated detection performance under occluded point clouds by removing points invisible from a given rendering angle and projecting the remaining ones. In some cases, abnormal regions were completely hidden, effectively converting anomalies into normal instances. To reduce this effect, we selected angles that preserved visibility of anomalous regions. Table XIII shows that PointAD+ exhibits performance degradation when tested on the occluded point clouds. We attribute this degradation to two factors: first, occluded point clouds may lose part of the anomaly semantics, making anomalies harder to recognize; second, occlusion can create artificial sinkholes on the surface, which may cause PointAD+ to mistakenly identify these regions as hole anomalies. Although PointAD+ outperforms PointAD, its performance gain on the occluded dataset is smaller than on the original dataset. We analyze that PointAD+ leverages intrinsic point relations. When points in certain regions are completely removed, these relations are disrupted and hinder geometric modeling.

### v5

@@ 5.1 Ablation on modules and key hyperparameters
TABLE VIII: Analysis on aggregation strategy $\ell$.
$\ell$ Point detection Multimodal detection
Local Global Local Global
Mean (96.9, 88.7) (82.5, 94.8) (98.2, 93.1) (87.7, 96.6)
Max (97.0, 88.3) (82.5, 94.4) (97.9, 93.5) (87.8, 96.4)
(a) The effect of $\alpha$ to the segmentation performance.
@@ E Point detection Multimodal detection
Ablation on aggregation strategy. By default, we adopt the mean aggregation strategy to obtain explicit point representations. We also investigate two alternative strategies: max strategy, which selects the maximum pixel representation across different views as the corresponding explicit point representation; network-parametric strategy, which employs an MLP followed by a softmax layer to compute aggregation weights of the pixel representations. As shown in Table VIII, PointAD+ is not restricted to a specific aggregation scheme but generalizes across different strategies. We adopt the mean strategy as it provides a favorable trade-off between performance and computational overhead.
@@
Ablation on rendering quality. In our experiments, we relied on the Open3D Library to generate point cloud renderings, though it does not provide an API for explicitly controlling rendering quality. Here, we apply a Gaussian blur to the 2D renderings to simulate the noise level. We evaluated PointAD on MVTec3D-AD under blur levels $\sigma\in\{1,5,9\}$. As summarized in Table XIV, the detection performance of PointAD+ decreases as the rendering quality degrades (i.e., with larger $\sigma$ values). Nevertheless, even under heavy blur ($\sigma=9$), the model maintains acceptable accuracy and consistently outperforms baselines that rely on high-quality renderings.
Ablation on point occlusion. We next evaluated detection performance under occluded point clouds by removing points invisible from a given rendering angle and projecting the remaining ones. In some cases, abnormal regions were completely hidden, converting anomalies into normal instances. To avoid this effect, we selected angles that preserved visibility of anomalous regions. Table XIII shows that PointAD+ exhibits performance degradation when tested on the occluded point clouds. We attribute this degradation to two factors: first, occluded point clouds may lose part of the anomaly semantics, making anomalies harder to recognize; second, occlusion can create artificial sinkholes on the surface, which may cause PointAD+ to mistakenly identify these regions as hole anomalies. Although PointAD+ outperforms PointAD, its performance gain on the occluded dataset is smaller than on the original dataset. We analyze that PointAD+ leverages intrinsic point relations. When points in certain regions are completely removed, these relations are disrupted and hinder geometric modeling.
@@ 6 Conclusion
Refer to caption Qihang Zhou received the B.S. degree from China University of Geosciences, Wuhan, China, in 2020. He is currently pursuing the Ph.D. degree in control science and engineering with the College of Control Science and Engineering, Zhejiang University, Hangzhou, China. His research interests include Vision-language models, anomaly detection, medical analysis, and ZS learning.
Refer to caption Shibo He (Senior Member, IEEE) received the Ph.D. degree in control science and engineering from Zhejiang University, Hangzhou, China, in 2012. He is currently a Professor with Zhejiang University. He was an Associate Research Scientist from March 2014 to May 2014, and a Postdoctoral Scholar from May 2012 to February 2014, with Arizona State University, Tempe, AZ, USA. From November 2010 to November 2011, he was a Visiting Scholar with the University of Waterloo, Waterloo, ON, Canada. His research interests include Internet of Things, crowdsensing, big data analysis, etc. Prof. He serves on the editorial board for the IEEE TRANSACTIONS ON VEHICULAR TECHNOLOGY, Springer Peer-to-Peer Networking and Application and KSII Transactions on Internet and Information Systems, and is a Guest Editor for Elsevier Computer Communications and Hindawi International Journal of Distributed Sensor Networks. He was a General Co-chair for iSCI 2022, Symposium Co-Chair for the IEEE/CIC ICCC 2022, IEEE GlobeCom 2020 and the IEEE ICC 2017, TPC Co-Chair for i-Span 2018, a Finance and Registration chair for ACM MobiHoc 2015, a TPC Co-Chair for the IEEE ScalCom 2014, a TPC Vice Co-Chair for ANT 2013 and 2014, a Track Co-Chair for the Pervasive Algorithms, Protocols, and Networks of EUSPN 2013, a Web Co-Chair for the IEEE MASS 2013, and a Publicity Co-Chair of IEEE WiSARN 2010, and FCN 2014.
Refer to caption Jiangtao Yan received the B.S. degree from Huazhong University of Science and Technology, Wuhan, China, in 2022. He is currently pursuing the M.S. degree in control science and engineering with the College of Control Science and Engineering, Zhejiang University, Hangzhou, China. His research interests include anomaly detection and Vision-language models.
Refer to caption Wenchao Meng (Senior Member, IEEE) received the Ph.D. degree in control science and engineering from Zhejiang University, Hangzhou, China, in 2015, where he is currently with the College of Control Science and Engineering. His current research interests include adaptive intelligent control, cyber–physical systems, renewable energy systems, and smart grids.
Refer to caption Jiming Chen (Fellow, IEEE) received the PhD degree in control science and engineering from Zhejiang University, Hangzhou, China, in 2005. He is currently a professor with the Department of Control Science and Engineering, the vice dean of the Faculty of Information Technology, Zhejiang University. His research interests include IoT, networked control, wireless networks. He serves on the editorial boards of multiple IEEE Transactions, and the general co-chairs for IEEE RTCSA’19, IEEE Datacom’19 and IEEE PST’20. He was a recipient of the 7th IEEE ComSoc Asia/Pacific Outstanding Paper Award, the JSPS Invitation Fellowship, and the IEEE ComSoc AP Outstanding Young Researcher Award. He is an IEEE VTS distinguished lecturer. He is a fellow of the CAA.